OpenStudy (anonymous):

Solve by factoring. Mark all solutions. 2x^2 + 3x - 14 = 0

hero (hero):

2x^2+3x-14 = 0 2x^2 + 7x - 4x - 14 = 0 x(2x+7)-2(2x+7) = 0 (2x+7)(x-2) = 0 x = 2 x = -7/2